Use Weierstrass’ M-Test to prove that the series


Infinity


∑ n^2x^n


n=1


converges uniformly in the interval [0,1/5]

1
Expert's answer
2022-03-14T19:16:53-0400

We have

"\left| n^2x^n \right|\leqslant \frac{n^2}{5^n}"

The series "\sum{\frac{n^2}{5^n}}" converges by D’Alembert’s test since

"\underset{n\rightarrow \infty}{\lim}\frac{\frac{n^2}{5^n}}{\frac{\left( n+1 \right) ^2}{5^{n+1}}}=5>1"

Then by Weierstrass’ M-Test the series "\sum{n^2x^n}" converges uniformly.